ADDING AND SUBBTRACTING RADICALS
HOW TO:
16 (square root)10 + 2 (square root)10
(SQUARE ROOT)5 - 3 (SQUARE ROOT)5
(SQUARE ROOT)3 (SQUARE ROOT)2 + 2 (SQUARE ROOT)3)

Answer by edjones(8007)

16sqrt(10)+2sqrt(10)=18sqrt(10)
.
sqrt(5)-3sqrt(5)=-2sqrt(5)
.
sqrt(3)*sqrt(2)=sqrt(3*2)=sqrt(6)
sqrt(6)+2sqrt(3) Ans.
.
Ed